# Co-Occ: Coupling Explicit Feature Fusion with Volume Rendering Regularization for Multi-Modal 3D Semantic Occupancy Prediction

### v2

@@ Abstract
{textblock*}
(2cm,0.5cm)This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ VI-A1 The Voxel Interaction of LiDAR and Camera
VI-A2 Other Implementation Details
VI-B Evaluation Metrics
$\rm{IoU}=\frac{TP_{o}}{TP_{o}+FP_{o}+FN_{o}},$ (11)
$\rm{mIoU}=\frac{1}{N_{c}}\sum_{c=1}^{N_{c}}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}},$ (12)
@@ VI-C More Results
More Visualization Results on NuScenes Dataset. Due to space constraints, we have included additional visual results in Table VII. These results demonstrate that our method has more precise details compared to the camera-only method [16], while also achieving greater consistency than other LiDAR-camera fusion-based methods [12]. Furthermore, the effectiveness of our methods in challenging conditions is validated through the video demo.

### v3

@@ Abstract
Index Terms:
Deep learning for visual perception, sensor fusion, semantic scene understanding
@@ II RELATED WORK
Scene Understanding with Volume Rendering. Recently, 3D perception approaches inspired by Neural Radiance Field (NeRF) [15] have employed volume rendering to generate 3D scene representations or estimate 3D geometry information from multi-view images. Previous studies, such as [32, 33], have focused on generating implicit scene representations through per-scene optimization using multi-view images. However, this approach can pose challenges when it comes to generalizing the results to different scenes. Approaches like [25, 26, 34] utilize NeRF to directly obtain 3D semantic occupancy predictions or acquire the rendered 2D semantic to supervise the semantic information. Unlike previous methods, we propose an implicit volume rendering-based regularization to enforce regularization on the fusion of LiDAR and camera data, based on the RGB and depth, to further enhance the unified volumetric representation.
@@ VI-A1 The Voxel Interaction of LiDAR and Camera
VI-A2 Other Implementation Details
VI-B Evaluation Metrics
$\rm{IoU}=\frac{TP_{o}}{TP_{o}+FP_{o}+FN_{o}},$ (11)
$\rm{mIoU}=\frac{1}{N_{c}}\sum_{c=1}^{N_{c}}\frac{TP_{c}}{TP_{c}+FP_{c}+FN_{c}},$ (12)
@@ VI-C More Results
More Visualization Results on NuScenes Dataset. Due to space constraints, we have included additional visual results in Fig. 9. These results demonstrate that our method has more precise details compared to the camera-only method [16], while also achieving greater consistency than other LiDAR-camera fusion-based methods [12]. Furthermore, the effectiveness of our methods in challenging conditions is validated through the video demo.